Cast Iron Grills, Inc., manufactures premium gas barbecue grills. The company uses a periodic inventory system and the LIFO cost method for its grill inventory. Cast Iron's December 31, 2018, fiscal year-end inventory consisted of the following (listed in chronological order of acquisition):

Units Unit Cost
8,800 $ 600
5,900 700
9,800 800


The replacement cost of the grills throughout 2019 was $900. Cast Iron sold 46,000 grills during 2019. The company's selling price is set at 200% of the current replacement cost.

Required:
1. & 2. Compute the gross profit (sales minus cost of goods sold) and the gross profit ratio for 2019 under two different assumptions. First, that Cast Iron purchased 47,000 units and, second, that Cast Iron purchased 24,500 units during the year.
4. Compute the gross profit (sales minus cost of goods sold) and the gross profit ratio for 2019 assuming that Cast Iron purchased 47,000 units (as per the first assumption) and 24,500 units (as per the second assumption) during the year and uses the FIFO inventory cost method rather than the LIFO method.
Answer #1

1&2. Under LIFO method, it is assumed that inventory purchased last is sold first.

Sales Units = 46,000

Selling Price = 200%*Replacement Cost = 200%*900 = $1,800

Cast Iron purchased 47,000 units

Sales 46,000*1800

82,800,000

Less: Cost of goods sold 46,000*900

41,400,000

Gross Profit

$41,400,000

Gross Profit Ratio = Gross Profit/Sales

= 50%

Cast Iron purchased 24,500 units during the year.

Sales 46,000*1800

82,800,000

Less: Cost of goods sold (24,500*900+9,800*800+5,900*700+5,800*600)

37,500,000

Gross Profit

$45,300,000

Gross Profit Ratio = 45,300,000/82,800,000

= 54.71%

4.FIFO method assumes that inventory purchased first will be sold first.

Cast Iron purchased 47,000 units:

Sales 46,000*1800

82,800,000

Less: Cost of goods sold (8,800*600+5,900*700+9,800*800+21,500*900)

36,600,000

Gross Profit

$46,200,000

Gross Profit Ratio = 46,200,000/82,800,000

= 55.80%

Cast Iron purchased 24,500 units during the year:

Sales 46,000*1800

82,800,000

Less: Cost of goods sold (8,800*600+5,900*700+9,800*800+21,500*900)

36,600,000

Gross Profit

$46,200,000

Gross Profit Ratio = 46,200,000/82,800,000

= 55.80%